OpenStudy (anonymous):

A crowbar is used as a lever. The effort force of 40 newtons moves 3 meters. The resistance force of 54 newtons moves 2 meters. What is the efficiency of the lever?

OpenStudy (anonymous):

work done was (40N)(3m) = 120 N-m work achieved was (54N)(2 m)= 108 N-m efficiency = achieved (output)/ effort (input) = 108/120